OpenStudy (anonymous):

Simplify (1/25)^-3/2

OpenStudy (anonymous):

125

OpenStudy (jdoe0001):

\(\bf \large { \left(\cfrac{1}{25}\right)^{-\frac{3}{2}}\implies \left(\cfrac{25}{1}\right)^{+\frac{3}{2}}\implies 25^{\frac{3}{2}} \\ \quad \\ recall\implies a^{\frac{{\color{blue} n}}{{\color{red} m}}} = \sqrt[{\color{red} m}]{a^{\color{blue} n}} \qquad \qquad \sqrt[{\color{red} m}]{a^{\color{blue} n}}=a^{\frac{{\color{blue} n}}{{\color{red} m}}}\quad thus \\ \quad \\ 25^{\frac{{\color{blue}{ 3}}}{{\color{red}{ 2}}}}\implies ? }\)